OpenStudy (waheguru):

PLZZ HELP ME a line has a slope of -2 and passes through -5,4 make and equation and change it to standard form PLZZ HELP ME a line has a slope of -2 and passes through -5,4 make and equation and change it to standard form @Mathematics

OpenStudy (waheguru):

help me guys

OpenStudy (anonymous):

what is standard form?

OpenStudy (anonymous):

y-b=m(x-a) in this case m = -2, b = 4, a = -5

OpenStudy (waheguru):

ok just give me the eqation of the line

OpenStudy (anonymous):

\[y - 4 = -2(x + 5)\]

OpenStudy (anonymous):

\[y = -2x - 6\]

OpenStudy (waheguru):

now change the formula to ay then by then c = 0

OpenStudy (waheguru):

helpp

myininaya (myininaya):

what?

OpenStudy (anonymous):

\[2x + y + 6 = 0\]

OpenStudy (waheguru):

a line has a slope of -2 and passes through -5,4 make and equation and change it to standard form

myininaya (myininaya):

oh the then's are plus signs lol

myininaya (myininaya):

and thats ax not ay got it

OpenStudy (waheguru):

yes adaddadadadad or what eva thx bro

OpenStudy (anonymous):

\begin{array}l\color{#FF0000}{\text{N}}\color{#FF0000}{\text{ }}\color{#FF7F00}{\text{o}}\color{#FF7F00}{\text{ }}\color{#FFE600}{\text{ }}\color{#FFE600}{\text{ }}\color{#FFE600}{\text{P}}\color{#FFE600}{\text{ }}\color{#00FF00}{\text{r}}\color{#00FF00}{\text{ }}\color{#0000FF}{\text{o}}\color{#0000FF}{\text{ }}\color{#6600FF}{\text{b}}\color{#6600FF}{\text{ }}\color{#8B00FF}{\text{l}}\color{#8B00FF}{\text{ }}\color{#FF0000}{\text{e}}\color{#FF0000}{\text{ }}\color{#FF7F00}{\text{m}}\color{#FF7F00}{\text{ }}\color{#FFE600}{\text{!}}\color{#FFE600}{\text{ }}\color{#00FF00}{\text{ }}\color{#00FF00}{\text{ }}\color{#00FF00}{\text{:}}\color{#00FF00}{\text{ }}\color{#0000FF}{\text{-}}\color{#0000FF}{\text{ }}\color{#6600FF}{\text{D}}\color{#6600FF}{\text{ }}\end{array}